On the 31/August Brain ltd has opening inventory of 20,000 goods. During September requisitions were raised for 23,000 valued at $12,000

Opening Cash was $7,000

Brain purchase 4,000 Goods at $2 and for 8,000 on Credit valued at $2 each from Law ltd

20,000 Goods were sold on credit totaling at $60,000 and 3,000 Cash Sales valued $9,000.

Of the 20,000 sold on credit the customer returned 1,000.

a)    Extract Brain Ltd retail’s profit & loss statement & balance sheet elements

1
Expert's answer
2021-06-02T12:34:34-0400

Solution:

a.). Brain Ltd retail’s profit & loss statement and balance sheet:

Profit and loss elements include the following:

Sales = 60,000 + 9,000 = 69,000

Less Sales returns = 522

Opening stock = 10,435

Add purchases = 24,000

Less closing stock = 6,261

Cost of sales = 28,174

 

Gross profit = 40,304

 

 

Balance sheet elements:

Current assets:

Cash = 7,000

Inventory = 6,261

 

Current liabilities:

Accounts payable = 16,000 + 60,000 = 76,000